Notes from 2007
This area contains slides of talks and reports
on work in progress on the End Cap Muon Alignment and software. If you
are not one of our collaborators, you are not entitled to review this
material. It contains preliminary work which can be inaccurate, and
only members of our collaboration are in a position to determine
- 08Jan2007 Found some events in
the sample with both lasers on at once. Status of new readout.
- 17Jan2007 Report on DCOPS
- 19Jan2007 I now have the
histograms for the raw data. Beware! For some of the events both lasers
were on. I also have a display of some histograms as a function of time.
- 22Jan2007 I found some more
events with both lasers on, and removed them from the lists.
- 24Jan2007 Pseudo-rphi deviations.
- 25Jan2007 Sorted out the SLM1 data
by magnetic field strength and repeated the pseudo-rphi measurement.
- 26Jan2007 Success! I can compare
Z displacements with photogrammetry measurements!
- 30Jan2007 Cleanup helps with
resolution. Look at CCD1 and CCD3 displacements in SLM1 (CCD1 is rather
- 01Feb2007 Biweekly status report.
CCD3 vs photogrammetry in ME+2/SLM1
- 02Feb2007 Does the beam go out of
range in the CCD1/CCD3 direction in the Transfer Lines? Probably not. Added
Handler's fit information into the data area,
MTCC1 information, and profile-free datafiles for ME+2/SLM2 and 3.
- 05Feb2007 Cleaner look (than
last entry) at MTCC1 Transfer Line5.
- 06Feb2007 Further look at MTCC1
transfer lines. There is evidence of a permanent change in direction in
some of the lasers, and not others. Tipping on the mount?
- 07Feb2007 First look at ME+2/SLM2
- 20Feb2007 Bob Handler looks at beam position vs field
- 22Feb2007 Look at ME+2/SLM2, include
photogrammetry for Z.
- 27Feb2007 Pictures of profiles
- 01Mar2007 Tried to put in calibrations.
Doesn't help much. Why?
- 23Mar2007 PickOne modified to
fit actual slice test cabling.
- 26Mar2007 SQL Command works now,
can write DCOPS_Configuration BLOB and read it back.
- 27Mar2007 DIM DNS node is pccmsalign01. There's a limit to the maximum number of open cursors which I hit if
something goes wrong.
- 16Apr2007 Fixed the bug that kept
the profiles from being written when the fit failed. Can't test until
cmsatsminus comes back, though.
- 19Apr2007 Recreated calibrated
versions of mep2slmx.stripped files.
- 23Apr2007 Created documentation
for new database tables.
- 26Apr2007 Bug fix in DCOPS_READOUT2
- 27Apr2007 ME-2 profiles requested,
provided. I modified the PickOne tool to create gif files.
- 18May2007 Status of calibration database.
- 22May2007 Evidence of a bug?
- 29May2007 Studying bug.
- 18Jun2007 CMS week reports from me
- 19Jun2007 DCOPS scans to align lasers.
- 21Jun2007 Software work and DCOPS
- 23Jun2007 Debugging the DAQ. Group
init bug found.
- 05Jul2007 Progress with the DAQ:
major group ID bug fixed. Calibration chain established: given a time,
find the DCOPS for an SLM nearest to that time, read them into appropriate
objects, and calibrate them.
- 11Jul2007 First pass at a calibrated DCOPS
data table. Look at quality measures. Presentation for 12-July
- 16Jul2007 Day vs night for ME-2/SLM2
- 17Jul2007 Oracle timing issues with
- 19Jul2007 ME-3/SLM1 profiles report,
and new DAQ software layout.
- 25Jul2007 ME-2/SLM1 profiles report,
changes to DAQ to handle DIM issues.
- 30Jul2007 Why so many zombies?
- 01Aug2007 Reports on DAQ status,
cover utility, and problems with Oracle and bug fixes. Animated gifs showing
effects of covers.
- 09Aug2007 Progress report on
Phoenix and TTree I/O.
- 27Aug2007 Notching at PT4 result.
I see improvement.
- 29Aug2007 Can only reap processes
that are children of the calling process, not siblings. Phoenix process
works. To first order, DAQ is done.
- 07Sep2007 Profiles retrieved from
Oracle put in root files, displayed in Madison. Oddity found.
- 10Sep2007 More profiles.
- 11Sep2007 Report on DCOPS DAQ. It
is finished, aside from scripting and documentation to make life easy for the
DCS installation folks. Profiles from MTCC2 are now in the data area.
- 14Sep2007 Eartly's calibration
constants are in DCOPS_CALIB.
- 27Sep2007 Status of work on
final DAQ system. Comments on profiles.
- 28Sep2007 Progress in UXC. Single read
works: Lasers 1 and 3 in ME+1 don't; nor do the lasers for ME+3. They go about
halfway across in ME+2, and less than that in ME+4.
- 02Oct2007 ME-1/2/3/4 lights off.
- 03Oct2007 ME-1/2/3 PickOne
- 05Oct2007 Non-linearity? Oleg's ME-4
- 08Oct2007 ME-1/2/3/4 data
- 10Oct2007 Bob has some data with
rather odd features. Also we have full SLM lines for ME-4/-3/-2/-1 and ME+2/+1
thanks to Oleg and Vladimir's efforts. Unfortunately you can see some shadowing.
- 13Oct2007 Status of DCOPS calibration
for COCOA input. Looks good, but needs final testing. Also collecting summary tables
for SLM lines. Some analysis.
- 20Oct2007 Calibration doesn't help reduce (1)-(3) spread.
- 24Oct2007 With/without straight line shimming
- 25Oct2007 Profiles for ME+3/ME+4, missing from 10-Oct set
(not very good). Crosstalk
sometimes causes problems for sensors with low signal.
- 26Oct2007 Results and
conclusions from studying reflections in ME+2.
- 30Oct2007 Paint test.
- 31Oct2007 Report on proposals to
cleanup DCOPS profiles
- 14Nov2007 What variation do we expect
in Rphi when the field goes on.
- 28Nov2007 Recent DCOPS data is available.
- 29Nov2007 Running program to retrieve MTCC data in COCOA format.
- 30Nov2007 MTCC2 DCOPS data is available.
- 07Dec2007 Dec CMS Meeting report.
Cocoa can read recent data from root file.
- 11Dec2007 Profiles (partly obscured)
These slides are presented here for the use of CMS collaborators; both
for quick communication during phone conferences and for archival
purposes. They refer to internal development work, and are not
intended for publication to the general public. As of this date none
of our collaborators are blind, and since none but our
collaborators are likely to benefit from reading them, I see no
purpose in making the postscript files accessible to the blind.
if you have trouble accessing the information on this page.